OpenStudy (anonymous):

Haven't been able to get this one please help! Simplify the expression cos^2 x -------- 1 + sin x A) 1 - sin x B) tan^2 x C) sec^2 x D) 1 E) 1 + cos x

OpenStudy (anonymous):

\[\frac{cos^2(x)}{1+sin(x)}=\frac{1-sin^2(x)}{1+sin(x)}=\frac{(1-sin(x))(1+sin(x))}{1+sin(x)}\]

OpenStudy (anonymous):

\[=1-sin(x)\]